Two objects attract each other with a gravitational force of magnitude 1.00 x 10^(-8)N when separated by 20.0 cm. If the total mass of the 2 objects is 5.00 kg, what is the mass of each?
1
Expert's answer
2021-02-22T10:29:16-0500

Let the first mass be "m_1=m". Then, the second one will be "m_2=5.0\ kg -m."

Let’s write the Law of Universal Gravitation:


"F=\dfrac{Gm_1m_2}{r^2}."

Let's substitute the numbers:


"1.0\cdot10^{-8} = \dfrac{6.67\cdot10^{-11}\cdot(5.0m-m^2)}{(0.2)^2},""\dfrac{1.0\cdot10^{-8}\cdot(0.2)^2}{6.67\cdot10^{-11}}=5.0m-m^2,""m^2-5m+6=0."

This uadratic eqqquation has two roots:


"m_1=\dfrac{-b+\sqrt{b^2-4ac}}{2a}=\dfrac{5+\sqrt{5^2-4\cdot1\cdot6}}{2\cdot1}=3\ kg,""m_2=\dfrac{-b-\sqrt{b^2-4ac}}{2a}=\dfrac{5-\sqrt{5^2-4\cdot1\cdot6}}{2\cdot1}=2\ kg."

Answer:

"m_1=3\ kg, m_2=2\ kg."